20201004 LBTB Science

01:22 David is collimating

01:30 OSCO running
  • ao telemetry
    • 20201004_013115.log
    • 20201004_013358.log
    • LUCI1 ended at 55, started at 35

01:42 First science object
  • scipt sent 12.0 mag AO Ref, but very dim on Acquistion camera
  • choose new AO Ref 1' away, 13.66 mag
  • Loop Closed
  • did not turn on GOPT. Ratio = 0.04. too low
  • 1.2" on DIMM

02:15 Loop Open for unknown reason
  • Closed the loop manually from the FLAO GUI and set gains,
  • Used the "Power" Maneuver to get going again (0,0 offset on LUCI telescope service GUI)

02:26 collect second set to data on same field. No new preset

02:40 New Preset
  • bin 4
  • camera lens will not converge. X at 100
  • try preset with delta pa of 90

03:00 camera lens converge
  • bin 4
  • TN 033041 no GOPT. value 0.07

03:40 Two AO Ref stars too dim. I could barely see them on about 1.3 of the Acquisition image frames

03:55 Done with AO for tonight

-- DougMiller - 04 Oct 2020
Topic revision: r2 - 04 Oct 2020, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback